Please help me solve this problem.
The basal energy requirement B is the number of calories that a person needs to maintain the life process. For a 28-year old female with a height of 160 centimeters and a weight of 45 kilograms (kg), B is 1300 calories. If her weight increases to 50kg, then B is 1365 calories. There is a linear equation that expresses B in terms of her weight w. Find the equation and find the basal energy requirements if her weight is 53.2kg.

Answer by fractalier(6550)

You are given 2 data points, like two x and y's...you find the equation of a line the same way...
Here our data points are (45, 1300) and (50, 1365). First find slope...
m = (1365 - 1300) / (50 - 45)
m = 65 / 5
m = 13
Now use the point slope form to get the equation...
y - 1300 = 13(x - 45)
y - 1300 = 13x - 585
y = 13x + 715
Now if her weight was 53.2, plug that in for x and calculate y...
y = 13(53.2) + 715 =
y = 1406.6 calories